# Compact and Selective Disclosure for Verifiable Credentials

### v1

@@ IV-B Protocol
This section describes how CSD-JWT enables compact and selective disclosure of claims within VCs using a positive ECC-based accumulator [22]. An identity owner $h$, also called holder, can selectively disclose a subset of claims $C$ among those included in a $vc$ issued by a trusted entity $i$, also known as the issuer. Beyond credentials, $i$ provides holders with the plain text of their claims $c_{j}\in\mathcal{C}$, along with the corresponding witnesses $w_{j}\in\mathcal{W}$. This information is used by $h$ to selectively disclose claims, empowering the holder with maximum control over their data meeting the SSI requirements.
@@ 8       $wvc\leftarrow wvc\cup\{(c_{j},w_{j})\}$;
Setup. The issuer securely generates the security parameter $1^{\lambda}$ and provides it to the $\mathsf{Setup}()$ function, producing the accumulator parameters $par$. This $par$ comprises the random generator of the ECC curve $\mathcal{G}$, and a key pair $(pk_{i},sk_{i})$. The private key $sk_{i}$ is necessary to accumulate claims to produce the accumulator value and to generate the witnesses, while $pk_{i}$ is used to verify their membership. This public key must be shared through a verifiable data registry such as blockchain, ensuring its authenticity and integrity. The issuer $i$ uses the $\mathsf{AccumulatorSetup}()$ function to initialize the accumulator obtaining the initial accumulator value $a$. This setup phase occurs only once, as the generated keys can support the management of multiple credentials.
Issuance of Verifiable Credential. The protocol begins with an identity owner $h$ requesting a $vc$ from a trusted entity $i$ (i.e., the issuer), which certifies a set of claims $\mathcal{C}=\{c_{1},...c_{n}\}$. Each claim is a key-value pair, composed of the name identifying a property and its value.
In CSD-JWT, instead of listing hashes for every $\mathcal{C}$, as for SD-JWT. CSD-JWT contains an accumulator value $a$, computed by accumulating all $c_{j}\in\mathcal{C}$, claims also include the holder and issuer identifier, respectively $DID_{h}$ and $DID_{i}$. The use of a unique identifier ensures that two distinct $vc$ containing the same $\mathcal{C}$ do not result in the same accumulator value $a$ and set of witnesses $\mathcal{W}$. The verifier uses $pk_{i}$, along with witness $W$, to verify whether $C$ is included in $a$. Algorithm 1 shows how VCs are issued in CSD-JWT.
The issuer executes the $\mathsf{AccumulateClaims}()$ function, providing the initial accumulator value $a$, $\mathcal{C}$, and $sk_{i}$, updating $a$. It is worth noting that, instead of accumulating $c_{j}$, we add $h(c_{j})$ to the accumulator to map it as a scalar value in the curve. Thus, it is highly unlikely that two distinct claims $c_{j}$ and $c_{k}$ produce the same digest, also ensuring that each witness corresponds to a unique claim. While not explicitly mentioned, in practice the issued $vc$ also contains all the necessary information that usually accompanies $vc$ such as the identifiers for $i$ and $h$ (i.e., credential structure or expiry date).
@@
[…]
Input: $vp$
Output: 00 or $1$
@@ 8             $res\leftarrow\mathsf{VerifyWitness}(a,c_{j},w_{j},pk_{i})$;
Verification of Verifiable Presentations. The verifier $v$ collects the $vp$ as JWT, including the accumulator value $a$, and the witness-value pairs $(w_{j},c_{j})\in WVC$ corresponding to the disclosed claims. Algorithm 2 shows how presentation are verified in CSD-JWT. To verify the claims presented by $h$, $v$ must be provided with the issuer’s and holder’s identifiers, previously included in $a$, which resolve to their respective DID Documents, containing their public keys, respectively, $pk_{i}$ and $pk_{h}$. The issuer’s DID Document is typically retrieved from a verifiable data registry, whereas the holder’s DID Document is usually provided directly by the holder.
Thus, $v$ executes the $\mathsf{VerifyPresentation}()$ function to verify the authenticity of claims included in the presentation, and to assess ownership over the credential. After verifying the holder’s signature through $pk_{h}$, the verifier proceeds by asserting that the disclosed claims $C$ are included in the accumulator value $a$ presented. Then, the verifier then executes the $\mathsf{VeirfyWitness}()$ function with $pk_{i}$ for each disclosed claims $c_{j}\in C$ and their corresponding witnesses $w_{j}\in\mathcal{W}$. Both the holder’s and the issuer’s identifiers are claims that must be validated in this process. As done by $i$ when computing $a$, before being combined with $w_{j}$ to prove inclusion in $a$, each $c_{j}$ is hashed with $h()$ using the same algorithm employed by $i$. If the combination of all the $h(c_{j})$ and $w_{j}$ matches $a$, the set of disclosed $C$ is authentic, and $v$ can trust the information shared.
@@ IV-C Discussion
Storage Requirement. For each claim in the VC, SD-JWT requires storing a 256-bit digest and a salt whose minimum length is 128 bits. In contrast, by leveraging the ECC-based accumulator proposed in [22] with a BN254 curve, CSD-JWT only requires storing a single 256-bit accumulator value and, for each claim to be included in the VC, a 256-bit witness. As the number of claims increases, our solution consumes less storage compared to SD-JWT. In Section VI, we comprehensively demonstrate that the reduction in storage space is about $(N-2)\times 256$ bits, where $N$ represents the number of claims included in the VC.
If the holder could perform runtime generation of $W$, we could further reduce the storage requirements of CSD-JWT. However, with the current state of cryptographic accumulators, the holder cannot generate $W$ while keeping the accumulated value at a fixed length [40]. This limitation is fundamentally tied to trapdoor-based accumulators, as enabling the holder to generate witnesses would require sharing the private key $sk_{i}$, allowing them to craft valid $W$ for claims $c_{j}\in C$ that have not been certified by any trusted party $i$. Therefore, to achieve runtime generation of constant-size $\mathcal{W}$, a trapdoor-less accumulator with fixed-length witnesses is needed, but such a construct does not currently exist.
Computing Requirement. CSD-JWT does not require intensive computations, making it suitable for constrained scenarios, such as hardware wallets or IoT devices, where computing capabilities may be limited. Using the accumulator, CSD-JWT does not impose any overhead on the holder side, as all the operations involving direct computation over this cryptographic structure (i.e., accumulating claims and generating corresponding witnesses) are performed on the issuer side.
On the other hand, generating VPs, as seen in SD-JWT and other approaches, must performed on the holder side since they are required to prove ownership of the disclosed claims. However, in CSD-JWT, the holder performs minimal computation, being only responsible for constructing and signing the JWT used for presentation, which contains pairs $(c_{j},w_{j})$ to disclose claims. Thus, the most resource-intensive operation for the holder is signing the JWT with their private key $sk_{h}$, which is the minimum requirement also for SD-JWT.
@@ V Security Analysis
This section examines the resilience CSD-JWT against the threats identified in the threat model described in Section III
@@ V-A Replay Attacks
Only the legitimate holder in control of a private key should be able to generate a valid $vp$. Without such guarantees, adversaries could impersonate credential holders by replaying their $vp$ or exploiting leaked data. CSD-JWT enforces this requirement by binding the presentation of claims to the holder’s DID. This ensures that credentials cannot be reused by third parties, even in cases of collusion or credential theft. We assume the unforgeability of asymmetric cryptographic mechanisms used by both the issuer and the holder, and the binding of DIDs to key pairs.
@@
A selective disclosure scheme is resilient against replay attacks if, for any probabilistic polynomial-time (PPT) adversary $\mathcal{A}$ who does not possess the holder’s private key $sk_{h}$, the probability that $\mathcal{A}$ can produce a $vp$ that is accepted by a verifier is negligible.
Let us consider an adversary $\mathcal{A}$ who obtains a valid credential bound to a legit holder $h$ by collusion or through theft. The adversary attempts to generate a valid presentation to impersonate $h$ to a verifier. Each credential includes the holder identifier $DID_{h}$. The adversary $\mathcal{A}$ cannot produce a valid witness corresponding to their identifier $DID_{\mathcal{A}}$ as this would require the issuer’s private key $sk_{i}$. Moreover, in CSD-JWT, each presentation $vp$ includes a cryptographic proof of possession of $sk_{h}$, in the form of a digital signature over a $vp$ containing a verifier-provided nonce $n$. Without $sk_{h}$, the adversary cannot generate a valid signature, due to the assumed unforgeability of the holder’s signature scheme. Even if $h$ colludes with $\mathcal{A}$ and provides a valid $vp$, the verifier’s use of a fresh nonce $n$ in each authentication session ensures that previously issued proofs cannot be replayed. Thus, $\mathcal{A}$ cannot reuse any presentation proof without interaction with $h$.
To further mitigate risks from these attacks, additional defenses should be enforced at the service provider level. These include implementing multi-factor authentication [41], monitoring usage patterns to detect suspicious activities [42], and using device attestation techniques to ensure that credentials are presented only from authorized devices [43].
@@ V-B Data Overcollection
Identity owners should provide only minimal information to verifiers such as service providers. Excessive collection of claims from a service provider could increase the risks for the holder, independently of its honesty. Malicious service providers may exploit data for financial gain or profiling purposes [44]. However, as seen for centralized models, honest servers could expose users to data breaches [45]. CSD-JWT overcomes this challenge by enabling $h$ to selectively disclose only the claims necessary to access a service, along with the corresponding witnesses. Thus, the service provider can verify the authenticity and integrity of the shared information, without seeing the full content of the credential.
Definition 3 (Data Minimization).
@@
A selective disclosure scheme satisfies data minimization if, during a verification session, the verifier $v$ learns only (i) the subset of claims $C\subseteq\mathcal{C}$ intentionally disclosed by the holder $h$; and (ii) auxiliary information that is independent of the remaining claims $\mathcal{C}\setminus C$ and their number.
During a credential presentation, the holder $h$ transmits: the disclosed claims $C$, cryptographic witnesses $W$, and an accumulator value $a$ committing to the full set $\mathcal{C}$ without revealing any additional information. The proof system used in CSD-JWT ensures that the witness $W$ is both sound (i.e., the claim truly belongs to the credential) and reveals nothing about $\mathcal{C}\setminus C$. The commitment $a$ is constructed using a cryptographic accumulator, which is collision-resistant and hiding and ensures that it is computationally infeasible to determine the number or content of accumulated values.
@@
This contrasts with schemes such as SD-JWT, where structural metadata (e.g., position or number of undisclosed claims) may be revealed during verification, potentially enabling side-channel inference attacks. For instance, consider a scenario where an individual holds a credential issued by a healthcare provider containing 20 claims related to their health condition. The holder wants to share only the health center and the last check-update with a fitness center to qualify for a health-based discount. By providing the entire credential for verification, $v$ can observe that many claims are not being disclosed. This observation can lead the verifier to speculate about the sensitive nature of the omitted information, even if those claims were not relevant to the service.
@@
CSD-JWT is resilient by design against this class of attacks because beyond the disclosed claims $C\subseteq\mathcal{C}$, the only additional information revealed are $a$ and $W$. This information does not leak anything about the full set of claims, specifically how many of them are contained in it.

### v2

@@ IV-B Protocol
This section describes how CSD-JWT enables compact and selective disclosure of claims within VCs. An identity owner or holder $h$ can selectively disclose a subset of claims $C$ among those included in a $vc$ issued by a trusted entity $i$, also known as the issuer. Beyond credentials, $i$ provides holders with the plain text of their claims $c_{j}\in\mathcal{C}$, along with the corresponding witnesses $w_{j}\in\mathcal{W}$. This information is used by $h$ to selectively disclose claims, empowering the holder with maximum control over their data, meeting the SSI requirements.
@@ 8  $wvc\leftarrow wvc\cup\{(c_{j},w_{j})\}$;
Setup. The issuer securely generates the security parameter $1^{\lambda}$ and provides it to the $\mathsf{Setup}()$ function, producing the accumulator parameters parpar. This parpar comprises, among other data required for the setup, a key pair $(pk_{i},sk_{i})$. The private key $sk_{i}$ is necessary to accumulate claims and produce the accumulator value, as well as to generate the witnesses, while $pk_{i}$ is used to verify their membership. This public key must be shared through a verifiable data registry, such as a blockchain, to ensure its authenticity and integrity. The issuer $i$ uses the $\mathsf{AccumulatorSetup}()$ function to initialize the accumulator, obtaining the initial accumulator value $a$. This setup phase occurs only once, as the generated keys can support the management of multiple credentials.
Issuance of Verifiable Credential. The protocol begins with an identity owner $h$ requesting a $vc$ from a trusted entity $i$ (i.e., the issuer), which certifies a set of claims $\mathcal{C}=\{c_{1},...,c_{n}\}$. Each claim is a key-value pair, composed of the name identifying a property and its value. CSD-JWT contains an accumulator value $a$, computed by accumulating all $c_{j}\in\mathcal{C}$, claims also include the holder and issuer identifier, respectively $DID_{h}$ and $DID_{i}$. The use of a unique identifier ensures that two distinct $vc$ containing the same $\mathcal{C}$ do not result in the same accumulator value $a$ and set of witnesses $\mathcal{W}$. The verifier uses $pk_{i}$, along with witnesses $W$, to verify whether claims $C$ are included in $a$. Algorithm 1 shows how VCs are issued in CSD-JWT.
The issuer executes the $\mathsf{AccumulateClaims}()$ function, providing the initial accumulator value $a$, $\mathcal{C}$, and $sk_{i}$, updating $a$. It is worth noting that, instead of accumulating $c_{j}$, we add $h(c_{j})$ to the accumulator to map it as a scalar value. Thus, it is highly unlikely that two distinct claims $c_{j}$ and $c_{k}$ produce the same digest, also ensuring that each witness corresponds to a unique claim. While not explicitly mentioned, in practice, the issued $vc$ also contains all the necessary information that usually accompanies $vc$, such as the underlying cryptographic mechanisms, the credential type, and its expiry date.
@@
Generation of Verifiable Presentations. By holding $a$ and the $WVC$, $h$ may reveal all claims or selectively disclose a subset $C\subseteq\mathcal{C}$ to any verifier $v$, such as service providers. To do this, $h$ executes the $\mathsf{GenerateVerifiablePresentation}()$ function with the accumulator value $a$, the claims to disclose $C$, the corresponding witnesses $W\subseteq\mathcal{W}$, and a nonce $n$ provided by $v$, generating a verifiable presentation $vp$ signed with their private key $sk_{h}$. This presentation is shared as a JWT. Specifically, for each $c_{j}$ to be disclosed, $vp$ contains a claim-witness pair ($c_{j},w_{j})$. The nonce $n$ is used to prevent replay attacks, where an adversary might intercept $vp$ and reuse it to spoof $h$.
Input: vpvp
Output: 0 or 11
@@ 8     $res\leftarrow\mathsf{VerifyWitness}(a,c_{j},w_{j},pk_{i})$;
Verification of Verifiable Presentations. The verifier $v$ collects the $vp$ as JWT, including the accumulator value $a$, and the witness-value pairs $(w_{j},c_{j})\in WVC$ corresponding to the disclosed claims. Algorithm 2 shows how presentations are verified in CSD-JWT. To verify the claims presented by $h$, $v$ must be provided with the issuer’s and holder’s identifiers previously included in $a$, which resolve to their respective DID Documents, containing their public keys, respectively, $pk_{i}$ and $pk_{h}$. The issuer’s DID Document is typically retrieved from a verifiable data registry, whereas the holder’s DID Document is usually provided directly by the holder.
Thus, $v$ executes the $\mathsf{VerifyPresentation}()$ function to verify the authenticity of claims included in the presentation, and to assess ownership over the credential. After verifying the holder’s signature through $pk_{h}$, the verifier proceeds by asserting that the disclosed claims $C$ are included in the presented accumulator value $a$. Then, the verifier executes the $\mathsf{VerifyWitness}()$ function with $pk_{i}$ for each disclosed claim $c_{j}\in C$ and their corresponding witnesses $w_{j}\in\mathcal{W}$. Both the holder’s and the issuer’s identifiers are claims that must be validated in this process. As done by $i$ when computing $a$, before being combined with $w_{j}$ to prove inclusion in $a$, each $c_{j}$ is hashed with $h()$ using the same algorithm employed by $i$. If the combination of all the $h(c_{j})$ and $w_{j}$ matches $a$, the set of disclosed $C$ is authentic, and $v$ can trust the information shared.
@@ IV-C Discussion
Storage Requirement. For each claim in the VC, SD-JWT requires storing a 256-bit digest and a salt whose minimum length is 128 bits. In contrast, as shown in Section VI, by implementing the ECC-based accumulator proposed in [vitto_biryukov] with a BN254 curve, CSD-JWT would only require storing a single 256-bit accumulator value and, for each claim to be included in the VC, a 256-bit witness. As the number of claims increases, our solution consumes less storage compared to SD-JWT. In Section VI, we comprehensively demonstrate that the reduction in storage space is about $(N-2)\times 256$ bits, where $N$ represents the number of claims included in the VC.
If the holder could perform runtime generation of $W$, we could further reduce the storage requirements of CSD-JWT. However, with the current state of cryptographic accumulators, the holder cannot generate $W$ while keeping the accumulated value at a fixed length [ghosh_2016]. This limitation is fundamentally tied to trapdoor-based accumulators, as enabling the holder to generate witnesses would require sharing the private key $sk_{i}$, allowing them to craft valid witnesses for claims that have not been certified by any trusted party $i$. Therefore, to achieve runtime generation of constant-size $\mathcal{W}$, a trapdoor-less accumulator with fixed-length witnesses is needed, but such a construct does not currently exist.
Computing Requirement. CSD-JWT does not require intensive computations, making it suitable for constrained scenarios, such as hardware wallets or IoT devices, where computing capabilities may be limited. Using the accumulator, CSD-JWT does not impose any overhead on the holder side, as all the operations involving direct computation over this cryptographic structure (i.e., accumulating claims and generating corresponding witnesses) are performed by the issuer.
On the other hand, generating VPs, as seen in SD-JWT and other approaches, must be performed by the holder as they are required to prove ownership of the disclosed claims. However, in CSD-JWT, the holder performs minimal computation, being only responsible for constructing and signing the JWT used for presentation, which contains the pairs $(c_{j},w_{j})$. Thus, the most resource-intensive operation for the holder is signing the JWT with their private key $sk_{h}$, which is the minimum requirement also for SD-JWT.
Revocation. Credential revocation refers to the process of invalidating a previously issued credential so that it can no longer be trusted or used by the identity holder. Revocation mechanisms apply to the entire credential and are orthogonal to selective disclosure techniques. For example, Revocation List 2020 [rl] maintains bitstrings where each bit corresponds to the status of a credential identifier, while [mazzoccaevoke] leverages cryptographic accumulators to track valid VCs. If accumulators are employed to support both revocation and selective disclosure, this dual-purpose design can yield substantial memory saving at the library level, further minimizing overhead on resource-constrained devices.
Moreover, it is worth noting that verifiability relies on the issuer’s digital signature; thus, any modification (e.g., removal or update of a single claim) requires re-issuance of the credential, a common limitation of selective disclosure mechanisms. In SD-JWT, this entails re-signing the list of salted hashes. In CSD-JWT, revocation is handled by updating the accumulator and recomputing witnesses for the remaining valid claims.
@@ V Security Analysis
This section analyzes the security guarantees of CSD-JWT against the threats identified in Section III, under standard cryptographic assumptions. We model adversaries as probabilistic polynomial-time (PPT) algorithms and express security properties in terms of negligible functions with respect to the security parameter $\lambda$.
@@ V-A Replay Attacks
To prevent impersonation, only the legitimate holder in possession of the private key $sk_{h}$ should be able to generate a valid $vp$. Without this guarantee, an adversary could replay previously captured $vp$ or exploit leaked data to impersonate the holder. CSD-JWT enforces replay protection by binding each presentation to the holder’s decentralized identifier $DID_{h}$ and requiring a fresh digital signature over a nonce $n$ provided by the verifier and included in the $vp$. This signature proves possession of $sk_{h}$ and prevents reuse of stale presentations. We assume that the holder’s signature scheme is existentially unforgeable under chosen message attacks (EUF-CMA) [euf], and that DIDs are securely bound to key pairs.
@@
A selective disclosure scheme is resilient against replay attacks if, for any PPT adversary $\mathcal{A}$ who does not possess the holder’s private key $sk_{h}$, the probability that $\mathcal{A}$ can produce a valid $vp_{\mathcal{A}}$ accepted by a verifier is negligible in $\lambda$:
$\Pr[\mathsf{Verify}(vp_{\mathcal{A}},pk_{h})=1]\leq\text{negl}(\lambda).$
Let $\mathcal{A}$ be an adversary who obtains a credential bound to a legitimate holder $h$ via collusion or theft. Each credential includes $DID_{h}$, and any valid presentation must include a signature over a verifier-provided nonce $n$ using $sk_{h}$. Without $sk_{h}$, $\mathcal{A}$ cannot generate a valid signature due to the assumed EUF-CMA security of the signature scheme. Even if $h$ colludes with $\mathcal{A}$ and provides a valid $vp$, the verifier’s challenge with a fresh nonce $n$ ensures that previously issued proofs cannot be replayed. Thus, $\mathcal{A}$ cannot reuse any presentation without active cooperation from $h$.
To further mitigate replay risks, additional defenses should be enforced at the service provider level. These include implementing multi-factor authentication [mfa], monitoring usage patterns to detect suspicious activities [candid], and using device attestation techniques to ensure that credentials are presented only from authorized devices [seda].
@@ V-B Data Overcollection
CSD-JWT enables selective disclosure by allowing the holder $h$ to reveal only a subset $C\subseteq\mathcal{C}$ of claims, along with their corresponding cryptographic witnesses $W$ and an accumulator value $a$ committing to the full set $\mathcal{C}$.
@@
Identity owners should provide only minimal information to verifiers, such as service providers. Excessive claim collection, regardless of the verifier’s honesty, can increase privacy risks for the holder. Malicious service providers may exploit the gathered knowledge either for financial gain or for profiling purposes [fbcambridge], while even honest providers may inadvertently expose users to data breaches [saleem2020sok]. CSD-JWT overcomes this challenge by enabling $h$ to selectively disclose only a subset $C\subseteq\mathcal{C}$ of claims. This enables the verifier to validate the authenticity and integrity of the shared information, without accessing the full credential.
Definition 3 (Soundness).
Let $\mathcal{C}$ be the set of claims committed to by the accumulator value $a$, and let $C\subseteq\mathcal{C}$ be the subset of claims disclosed by the holder $h$ during a credential presentation. A selective disclosure scheme satisfies soundness if, for any PPT adversary $\mathcal{A}$, the probability that $\mathcal{A}$ can produce a tuple $(c_{\mathcal{A}},w_{\mathcal{A}},a)$ such that:
  • $c_{\mathcal{A}}\notin C$ (i.e., at least one disclosed claim $c_{a}$ is not part of the original credential),
  • $w_{\mathcal{A}}$ is a valid witness for $c_{\mathcal{A}}$ with respect to $a$,
  • and $\mathsf{VerifyWitness}(a,c_{\mathcal{A}},w_{\mathcal{A}},pk_{i})=1$,
is negligible in the security parameter $\lambda$:
$\Pr\left[\begin{array}[]{c}\mathsf{VerifyWitness}(a,c_{\mathcal{A}},w_{\mathcal{A}},pk_{i})=1\\ \land\ c_{\mathcal{A}}\notin C\end{array}\right]\leq\text{negl}(\lambda).$
Definition 4 (Information Minimality).
@@
A selective disclosure scheme satisfies data minimization if, during a verification session, the verifier $v$ learns only:
  • the subset of claims $C\subseteq\mathcal{C}$ intentionally disclosed by the holder $h$, and
  • auxiliary information (e.g., witnesses $W$, accumulator $a$) that is computationally independent of $\mathcal{C}\setminus C$ and its cardinality $|\mathcal{C}\setminus C|$.
In CSD-JWT, during the presentation of a credential, the holder $h$ transmits the disclosed claims $C$, cryptographic witnesses $W$, and an accumulator value $a$ committing to the full set $\mathcal{C}$ without revealing any additional information. Crucially, the proof system guarantees that the verifier can validate the authenticity of the disclosed claims without learning anything about the undisclosed portion $\mathcal{C}\setminus C$. This is achieved through two key properties: (i) soundness, ensuring that each disclosed claim in $C$ is verifiably part of the committed set $\mathcal{C}$, and (ii) information minimality, meaning that $W$ and $a$ reveal no information about the content, structure, or size of $\mathcal{C}\setminus C$. As a result, CSD-JWT fully satisfies the data minimization definition by limiting verifier knowledge to only what is explicitly disclosed, while preserving the privacy of all other credential data.
It is worth noting that structural metadata, such as position or number of undisclosed claims, can itself leak sensitive information, as recently highlighted in the SD-JWT IETF draft [ietf-oauth-selective-disclosure-jwt-22]. To mitigate this, SD-JWT recommends adding decoy digests (i.e., mock claims) when a real claim is conditionally included. Specifically, if a claim is only present when a specific condition is met, the issuer should include a decoy digest when the condition is not met, to prevent revealing the absence of the claim.
@@
Let us consider a scenario where revealing the claim number can be exploited to enable manipulative practices or unfair decision-making. For instance, an individual holds a credential issued by a healthcare provider containing 20 claims related to their health condition. To qualify for a health-based discount, the holder wishes to share only two claims (i.e., health center and the last check-update with a fitness center). However, by providing the entire credential for verification, $v$ can observe that many claims are being withheld. This observation alone may lead the verifier to speculate about the sensitive nature of the undisclosed information, even if those claims were not relevant to the service.
@@
CSD-JWT is resilient by design against this class of attacks because beyond the disclosed claims $C\subseteq\mathcal{C}$, the only additional information revealed is $a$ and $W$. This information does not leak anything about the full set of claims, specifically how many of them are contained in it. In contrast, SD-JWT can achieve similar privacy guarantees through the use of decoys; however, this comes at the cost of increased credential size, making CSD-JWT even more efficient and impactful.